OpenStudy (anonymous):

Create a quadratic equation in standard form that can be factored. Part A. Write your equation in standard form. Use complete sentences to explain the benefits of writing your equation in standard form. Part B. Write your equation in factored form. Use complete sentences to explain the benefits of writing your equation in factored form. Part C. Write your equation in vertex form. Use complete sentences to explain the benefits of writing your equation in vertex form.

OpenStudy (anonymous):

Part D. Explain how all three forms can be used together to help you graph a quadratic function. Graph your function, and label the y-intercept, the x-intercepts, and the vertex.

OpenStudy (anonymous):

@triciaal @mathmale @mathstudent55 @zepdrix @sleepyjess @ganeshie8 @boldjon @pooja195 @ParthKohli @dan815 @nnesha

OpenStudy (anonymous):

@Michele_Laino @imqwerty @robtobey @.Sam. @welshfella @jagr2713 @ikram002p @inkyvoyd @ShadowLegendX @rvc @Awolflover1

OpenStudy (anonymous):

Can someone please help with this?

OpenStudy (anonymous):

First, make an equation that you can factorize

OpenStudy (anonymous):

Standard form is ax + by = c, correct?

OpenStudy (nincompoop):

Create a quadratic equation in standard form that can be factored. Part A. Write your equation in standard form. Use complete sentences to explain the benefits of writing your equation in standard form. First, what does a quadratic equation in standard form looks like?

OpenStudy (anonymous):

@nincompoop A quadratic equation is a 2nd degree equation, right?

OpenStudy (nincompoop):

yes

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

The Standard form of a quadratic equation is ax^2 + bx + c I'm assuming they want a quadratic in ax + by = c form, correct?

OpenStudy (nincompoop):

the first one

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Oh, ax^2 + bx + c? alright

OpenStudy (nincompoop):

EQUATION so do not forget the right side of the equation.

OpenStudy (nincompoop):

what you had was an EXPRESSION and not an equation.

OpenStudy (anonymous):

12x^2 + 4x + 8 = 0?

OpenStudy (nincompoop):

something like that

OpenStudy (anonymous):

"something like that" meaning something is wrong with my equation?

OpenStudy (nincompoop):

is that factorable ?

OpenStudy (anonymous):

well, 4 can definetely be factored out of it

OpenStudy (nincompoop):

go on and then?

OpenStudy (anonymous):

4(3x^2 + x + 2) = 0

OpenStudy (anonymous):

ermmm...

OpenStudy (anonymous):

hmm

OpenStudy (nincompoop):

so how would you factor the one inside the parenthesis in the left hand side?

OpenStudy (anonymous):

what one?

OpenStudy (nincompoop):

(3x^2 + x + 2)

OpenStudy (anonymous):

I'm getting really confused by your "one". Do you literally mean the number 1, or are you saying something else?

OpenStudy (nincompoop):

how do you factor the number 1? you cannot so it must be something else. How do you factor "the one INSIDE the parenthesis" on the left hand side. DO you have anything that is enclosed in the parenthesis? YES (3x^2 + x + 2) Now, factor it.

OpenStudy (anonymous):

a common factor between 3x^2 and x is x, but what would happen with the 2?

OpenStudy (nincompoop):

I don't know. You tell me.

OpenStudy (nincompoop):

you are the one that came up with the equation, not me.

OpenStudy (anonymous):

Is it possible to leave out a number from the equation when factoring?

OpenStudy (nincompoop):

meaning?

OpenStudy (anonymous):

as in, could I do 4((3x + 1)2) ?

OpenStudy (anonymous):

Wait, sorry that's wrong

OpenStudy (anonymous):

4(x(3x + 1)2)

OpenStudy (anonymous):

Like, is that possible in factoring, or do I have to include the 2 as well?

OpenStudy (nincompoop):

No. First, what is the number two doing there? is that suppose to be an exponent or what?

OpenStudy (anonymous):

4(3x^2 + x + 2) = 0

OpenStudy (anonymous):

That's what the number 2 is doing there I factored x out of 3x^2 + x, is that allowed or do I have to find a GCF for all 3 terms inside the parantheses

OpenStudy (nincompoop):

No. That is incorrect.

OpenStudy (anonymous):

alright

OpenStudy (nincompoop):

What is this? 4(x(3x + 1)2)

OpenStudy (anonymous):

So is there any way to factor 4(3x^2 + x + 2) = 0 anymore? Any technique that I don't know?

OpenStudy (anonymous):

4 ( 3x^2 + x + 2) = 0

OpenStudy (nincompoop):

You tell me! you're supposed to have learned this already.

OpenStudy (nincompoop):

You've been doing factoring for the past 3 weeks.

OpenStudy (anonymous):

4 ( 3x^2 + x + 2) = 0 GCF Between 3x^2 and x = x 4 ( x( 3x + 1 ) + 2) = 0

OpenStudy (anonymous):

That was my way of thinking when I got 4(x(3x + 1)2) Although i forgot the + sign before the 2

OpenStudy (nincompoop):

How about this. Start with something simpler.

OpenStudy (nincompoop):

change your equation to something you can actually factor properly.

OpenStudy (anonymous):

Fine @nincompoop In anycase, is 4(3x^2 + x + 2) fully factored?

OpenStudy (nincompoop):

Well that is the farthest that you can do. But that is not a good expression, because you can't really further factor whatever you have inside the parenthesis. Remember that you've been factoring quadratic equations to something like (x+a) (x+b)

OpenStudy (anonymous):

Is there some technique to know if a quadratic equation/expression is factorable in the first place?

OpenStudy (anonymous):

Or do I just have to guess and check?

OpenStudy (nincompoop):

YES. You reverse what you have been learning to factor.

OpenStudy (anonymous):

using FOIL, correct?

OpenStudy (nincompoop):

If you can factor \((x^2+8x+16) \) into \((x+4 ) (x+4) \) which gives you \((x+4)^2 \) then you can do the same thing in reverse way.

OpenStudy (anonymous):

(2x + 2)(4x + 3) First Inner Outer Last x * x = 8x^2. 2 * 4x = 8x. 2x * 3 = 6x. 2 + 3 = 6 8x^2 + 8x + 6x + 6

OpenStudy (anonymous):

Simplify: 8x^2 + 14x + 6

OpenStudy (anonymous):

Done

OpenStudy (anonymous):

Write your equation in standard form. Use complete sentences to explain the benefits of writing your equation in standard form.

OpenStudy (nincompoop):

it is an equation, so do not fail to forget the right-hand-side of the equation.

OpenStudy (anonymous):

So just to clarify, I just completed Part A correct? I'm confused. Does standard form mean ax + by = c or does standard form mean ax^2 + bx + c?

OpenStudy (anonymous):

8x^2 + 14x + 6 = 0

OpenStudy (nincompoop):

NO. This is the standard-form: \(ax^2+bx +c = 0 \)

OpenStudy (anonymous):

so what is ax + by = c?

OpenStudy (nincompoop):

8x^2 + 14x + 6 = 0 this is correct

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Write your equation in factored form. Use complete sentences to explain the benefits of writing your equation in factored form.

OpenStudy (anonymous):

factored form: (2x + 2)(4x + 3)

OpenStudy (nincompoop):

I don't know who taught you that ax + by = c is a quadratic equation First, your highest degree of polynomial is not two, it is one

OpenStudy (anonymous):

What are the benefits of having an equation in factored form?

OpenStudy (anonymous):

Oh I just looked it up

OpenStudy (nincompoop):

what is the benefits from what you can tell?

OpenStudy (anonymous):

Oh nevermind, that's standard form for a linear equation, not quadraic

OpenStudy (nincompoop):

exactly.

OpenStudy (anonymous):

It's neater, easier to find out what x equals? thats all i can think of

OpenStudy (nincompoop):

so what is the benefits of writing quadratic equations in standard form?

OpenStudy (anonymous):

easier to find out what x equals @nincompoop ?

OpenStudy (nincompoop):

how?

OpenStudy (anonymous):

sorry forgot the = 0 part again (2x + 2)(4x + 3) = 0

OpenStudy (nincompoop):

that is something you can tell in factored form

OpenStudy (anonymous):

(2x + 2) = 0 (4x + 3) = 0

OpenStudy (nincompoop):

It is not always easy to tell what the values of x are in \(ax^2+bx+c= 0 \)

OpenStudy (anonymous):

Then you can easily solve for x. Are there actually any benefits other than that for writing a quadratic equation/expression in factored form?

OpenStudy (anonymous):

@nincompoop ?

OpenStudy (nincompoop):

okay, let us test your answer but with a different equation. I will use what you originally had put 12x^2 + 4x + 8 = 0 Tell me what are the values of x in that equation?

OpenStudy (anonymous):

no idea

OpenStudy (nincompoop):

well first, you can't easily factor so you can't easily find your x values. So you can easily tell that it is not a perfect square trinomial. Perfect square trinomials are binomials multiplied by itself and an example is a binomial \(x+5\) when multiplied by itself becomes \((x+5)^2 \rightarrow (x+5)(x+5)\)

OpenStudy (anonymous):

To be honest, the only benefit I can think of is you can find the x-intercepts, or solutions, or x value, or whatever you want to call it easily if it is fully factored and the equation is equal to 0. Beyond that my mind is a blank

OpenStudy (anonymous):

Maybe that is the only benefit o_O

OpenStudy (anonymous):

Write your equation in vertex form. Use complete sentences to explain the benefits of writing your equation in vertex form.